# X-LeBench: A Benchmark for Extremely Long Egocentric Video Understanding

### v1

@@ Abstract
Long-form egocentric video understanding provides rich contextual information and unique insights into long-term human behaviors, holding significant potential for applications in embodied intelligence, long-term activity analysis, and personalized assistive technologies. However, existing benchmark datasets primarily focus on single, short-duration videos or moderately long videos up to dozens of minutes, leaving a substantial gap in evaluating extensive, ultra-long egocentric video recordings. To address this, we introduce X-LeBench, a novel benchmark dataset specifically crafted for evaluating tasks on extremely long egocentric video recordings. Leveraging the advanced text processing capabilities of large language models (LLMs), X-LeBench develops a life-logging simulation pipeline that produces realistic, coherent daily plans aligned with real-world video data. This approach enables the flexible integration of synthetic daily plans with real-world footage from Ego4D—a massive-scale egocentric video dataset covers a wide range of daily life scenarios—resulting in 432 simulated video life logs that mirror realistic daily activities in contextually rich scenarios. The video life-log durations span from 23 minutes to 16.4 hours. The evaluation of several baseline systems and multimodal large language models (MLLMs) reveals their poor performance across the board, highlighting the inherent challenges of long-form egocentric video understanding and underscoring the need for more advanced models.
@@ 1 Introduction
Understanding long-form egocentric videos captured from a first-person perspective over extended periods holds significant potential for advancing various domains such as embodied intelligence, long-term activity analysis, and personalized assistive technologies [30, 25, 29]. These videos provide rich contextual information and unique insights into human behaviors as they unfold naturally throughout the day. The ability to analyze such extensive recordings is key to developing more personalized agent systems that can construct long-term memory from users’ daily recordings, better understand human actions, anticipate needs, and interact more seamlessly in real-world settings [24, 17, 31].
@@
However, most existing datasets [37, 5, 18, 20, 43, 27], predominantly focus on short individual video clips captured from third-person views, failing to document detailed individual activities that make them insufficient for in-depth human-centered research. Egocentric video understanding has recently gained more attention, with several notable benchmarks and datasets being proposed [8, 9, 12, 50, 26, 13, 25, 10]. Ego4D [12] features massive-scale egocentric videos covering a wide range of daily activities, with video durations ranging from 5 seconds to 7 hours. Several subsequent research [26, 13, 16, 35, 6] based on Ego4D have also emerged. They fall short of capturing the full scope of long-term daily human activities, as they mainly focus on isolated clips or recordings. This limitation hinders the development and evaluation of models designed to process ultra-long video streams. In particular, it restricts the ability to challenge current long-form video understanding systems and models that construct long-term memory from video recordings and retrieve relevant information in response to user queries. Without extensive, continuous contextual data, evaluating the robustness of their performance becomes difficult.
@@
Creating benchmark datasets that span several hours of egocentric video is necessary but presents significant challenges. Data acquisition is a primary hurdle, recording dozens or even hundreds of hours-long egocentric videos is labor-intensive and raises privacy concerns, as it requires participants to wear recording devices for extended periods while going about their daily lives. Device limitations add to the complexity, as continuous video capture demands substantial storage capacity and consistent performance from recording equipment, which can affect data quality and reliability. Additionally, ensuring annotation quality in long-form videos is daunting. Manually annotating extensive recordings is time-consuming and can lead to annotator fatigue, compromising the detail and accuracy of the labels needed for training and evaluating complex models.
@@
To address these challenges, we introduce X-LeBench, a versatile and scalable benchmark dataset specifically designed for evaluating tasks on extremely long egocentric videos. X-LeBench features a life-logging simulation pipeline that synthesizes new data from existing short (a few seconds) or moderately long (a few hours) videos. It employs large language models (LLMs) to dynamically generate realistic daily plans aligned with real-world activities based on adjustable input settings. Specifically, this simulation integrates synthetic daily plans with actual footage from Ego4D, matches them in three core information aspects, and then dynamically and iteratively optimizes the simulation process based on the retrieved information, resulting in simulated video life logs that mirror daily activities in rich contexts with their duration extended to dozens of hours. A generated sample of the video life-log simulation is shown in Fig. 1, which consists of video recordings at different timestamps. Notably, X-LeBench offers a customizable and scalable design that allows to synthesis data of different lengths and content, facilitating future research of tailored needs.
@@
Our initial evaluations of baseline systems and multimodal LLM (MLLM) show consistently poor performance on X-LeBench. These results highlight the inherent difficulties of long-form egocentric video understanding and underscore the need for more advanced models. X-LeBench aims to facilitate research in this area by providing a comprehensive benchmark, encouraging the development of solutions that can effectively interpret and analyze ultra-long egocentric videos.
Our contributions are:
  • We present X-LeBench, the first benchmark dataset that encompasses ultra-long egocentric video recordings.
@@
  • We introduce a novel and customizable pipeline that simulates realistic, hours-long egocentric videos by integrating synthetic daily plans with real-world footage.
  • We conduct extensive evaluations of existing models on X-LeBench, revealing significant performance gaps and highlighting the challenges that future research must address.
@@ 2 Related Works
Dataset Avg. Duration (mins) Annotation Scheme Egocentric
MVBench [19] 0.27 Auto No
ActivityNet-QA [5] 1.85 Manual No
EgoSchema [26] 3 Auto&Manual Yes
EgoPlan-Bench2 [33] up to 5 Auto&Manual Yes
MovieChat-1K [36] 9.4 Manual No
MLVU [49] 12 Auto&Manual Partial
Video-MME (Short) [11] 1.37 Manual No
@@
HourVideo [6] 45.7 Auto&Manual Yes
InfiniBench [2] 52.59 Auto&Manual No
LVBench [42] 68.35 Manual No
Ours (Short) 142 Auto&Manual Yes
@@
2.1 Egocentric Video Understanding Benchmarks
Egocentric video understanding has received increasing attention in embodied intelligence, long-term activity analysis, and personalized assistive technologies [7, 14]. Ego4D [12] addresses the demand for extensive egocentric data by providing over 3,670 hours of egocentric videos across 74 locations. It enables significant advancements, but presents videos in isolated segments, lacking coherent, continuous recordings under real-life contexts.
Related works [4, 23, 31, 16] extend Ego4D, further exploring various applications of egocentric video understanding. EgoSchema [26] offers an egocentric video question-answering benchmark with over 5,000 curated multiple-choice pairs but focuses on 3-minute clips, insufficient for the evaluation of extended temporal understanding.
The Aria Everyday Activities dataset [25] offers multimodal egocentric data,
HourVideo [6] curates videos of dozens of minutes from Ego4D for evaluating hour-long video-language understanding. EgoPlan-Bench2 [33] assesses planning capabilities of MLLMs. However, these works either lack coherent, continuous daily-life recordings over long time spans or face data scale and diversity limitations. In contrast, our work addresses these gaps by building an ultra-long, continuous daily-life egocentric video dataset.
2.2 Long-form Video Understanding Benchmarks
The definition of “long” in video understanding varies across benchmarks, ranging from a few minutes to several hours, with most focusing on third-person videos. As shown in Tab. 1, EgoSchema [26] defines 3-minute videos as long using their proposed certificate length, Video-MME [11] classifies videos with a length of 30-60 minutes as long. HourVideo [6] and LVBench [42] specifically propose benchmarks for long video understanding, and define videos longer than 20 and 30 minutes as long videos respectively. MLVU [49] offers videos of diversified lengths with a 12-minute average length, providing comprehensive evaluation tasks for MLLMs’ long video understanding capabilities. InfiniBench [2] pushes the boundary with 50-minute videos and over 108,000 question-answer pairs, posing significant challenges for leading AI models.
Despite recent advances, these datasets are neither sufficiently long for evaluating ultra-long video processing nor focused on egocentric content, leaving a gap in addressing this emerging real-world application. Our work addresses this by redefining “long video” to include multi-hour, contextually consistent egocentric recordings with sufficient annotations to advance model and system development.
@@ 2.3 LLM-assisted Annotation Scheme
Traditional annotation processes require extensive human effort, but advances in LLMs and MLLMs [1, 44, 3, 38, 40, 21] have enabled their use in benchmarks. As shown in Tab. 1, a range of video understanding benchmarks leverage LLMs/MLLMs to automate the annotation process [26, 34, 32]. For instance, EgoSchema [26] generates question-answer pairs by querying an LLM with human-annotated narrations and tailored prompts, MVBench [19] uses LLMs to generate annotations for diverse video datasets.
In our work, the video duration extends to more than ten hours, making manual annotation virtually impractical and prone to fatigue-related errors. To address this, we also adopt an automatic and manual annotation scheme. First, we adapt Ego4D’s manual annotations to fit our task design. Then, we use LLMs to consolidate summaries of 5-minute video clips from Ego4D at various granularities, creating single-video, multi-video, and holistic-level summaries (detailed in Sec. 3.2.2).
@@ 3 X-LeBench
In real-world scenarios, life-logging with egocentric devices is constrained by hardware limitations and privacy concerns, preventing continuous recording. Thus, obtaining long-term life-logging video records is inevitably a combination of multiple video segments captured at different times. Building on this insight, we develop the life-logging simulation pipeline. By leveraging the text processing capabilities of LLMs and the extensive annotations of the Ego4D dataset, we generate simulated video life logs that realistically reflect daily activities and maintain contextual coherence, eventually culminate in the creation of our X-LeBench, detailed in the following sections.
@@ 3.1 Life-logging Simulation Pipeline
Stage 1 - Persona Generation. We dynamically generate personalized character profiles based on the input of different locations and the myers-briggs type indicators (MBTI) [28], generating basic background information including the character’s personality traits, lifestyle, hobbies and general daily routines. This information is the initial character’s settings to facilitate subsequent diverse and unique generations of detailed daily plans. We then generate a detailed daily activities agenda that matches the initial settings, and further use it to generate the daily plan chunks, where each chunk includes the activity plan of the character in different time periods. During this stage, we use GPT-4o [15] for character generation, setting 9 different locations based on the demographics of Ego4D dataset, and 16 MBTI types for each location, resulting in a total of 144 different persona profiles.
@@
Stage 2 - Video Information Extraction. The Ego4d dataset provides video content that spans hundreds of scenarios, allowing a glimpse into daily life around the world, and is labeled with dense video summaries for every 5-minute video clip. We construct our video selection library by carefully selecting 7852 videos from the dataset that are labeled with scenarios and clip-level summaries and not redacted. Then we extract three core aspects of video information from the library, including the time, scene, and the main activities presented in these videos. Specifically, we use Gemini 1.5 Pro [39] to analyze videos. To facilitate simulation with reasonable environment settings and lighting conditions, videos are categorized into different scene environments: indoor, outdoor, and mixed, as well as different time periods: daytime, nighttime, twilight, and uncertain. It is worth noting that the main scenes of some videos are indoors, making it difficult to recognize their time period information, for this kind of videos, we assign the special label “uncertain”. As for the extraction of content information, we directly extract the scenario information (i.e. main activity description) from Ego4D as the overview of the video content, and use Gemini to consolidate a series of 5-minute clip-level summaries into a 1-sentence summary of each video.
@@
Figure 3: Dataset statistics. (a): Video selection library consists of 7852 videos from Ego4D dataset, covering 135 scenarios, here shows the top 30. (b): X-LeBench has video life-log durations span from 23 minutes to 16.4 hours, here shows distribution of duration range across different data categories. (c) and (d) show the distribution of scene and time information categories of videos in video selection library, respectively. (e): Statistical information of duration lengths of different data categories, including minimum, mean, median and maximum values. (f): The distribution of number of videos with different occurrences.
@@
Stage 3 - New Dataset Generation. In this stage, we generate the new dataset by linking the generated daily plan chunks (from Stage 1) with the real video information (from Stage 2). For each chunk, we retrieve a video that aligns with its time, scene, and content, constructing the video life-log simulation. The selection process follows this rule:
$V={\underset{V\in\mathbf{M_{v}}}{{\arg\max}}\ S(\mathbf{M_{v}}(C_{t},C_{scene},C_{scenario}),C_{desc})}$ (1)
@@
where $V$ is the selected video for chunk $C$, determined by the alignment with $C$’s core information, which is time ($C_{t}$), scene ($C_{scene}$), scenario ($C_{scenario}$), and 1-sentence summary description ($C_{desc}$). $\mathbf{M_{v}(\cdot)}$ denotes the operation of matching videos based on the provided chunk information. $S(\cdot)$ represents the sentence similarity computation between $C_{desc}$ and 1-sentence summary set of matching videos. Specifically, $C_{t}$ and $C_{scene}$ are inferred from chunk timestamps and content, while $C_{scenario}$ is inferred using the dataset’s predefined scenario list (i.e. the dataset reference in Fig. 2).
@@
To efficiently identify the most suitable videos from the large-scale video selection library, we first apply a coarse-grained filtering step, i.e., the operation $\mathbf{M_{v}}(C_{t},C_{scene},C_{scenario})$, a rule-based matching to narrow down the candidate video set. Finally, to refine the selection, we calculate the sentence similarity between the 1-sentence summary of $C$ and the summaries in the candidate video set ($\mathbf{M_{v}}$), ensuring alignment across time, scene, and content dimensions.
@@
In addition, to enhance the contextual reasonableness and coherence of overall video life logs, we preferentially select videos recorded within the same country for each simulation. Furthermore, an iterative optimizing strategy is adopted, i.e., after each matching of a daily plan chunk, we update the life-log simulation memory with the information of selected video, which acts as the real daily activities record of the persona. This information, along with the remaining chunks to be simulated, is provided to the LLM. The LLM uses this context to reasonably adjust and update the subsequent daily plan chunks, ensuring that the overall simulation maintains logical coherence and contextual alignment.
Statistics. Our life-logging simulation pipeline integrates flexible parameterization settings. By specifying the location and MBTI-type information, we can get persona profiles of different styles, resulting in diverse simulations. Also, we can customize different nnumber of chuncks to retrieve and get video life logs of various video compositions. In this work, we set three different chunk numbers for simulation, namely 4, 9, and 15 (corresponding to short, medium, and long life-log categories), for each location-MBTI combination generated persona, thus obtaining 432 extremely long video life logs.
As summarized in Tab. 2, Fig. 3 (b) and (e), the duration statistics across categories show significant diversity: short data are mostly 1-2 hours, medium data are mostly 4-5 hours, and long data are mostly 6-8 hours, with the average duration of 2.37, 5.32 and 8.6 hours respectively.
Table 2: Statistics of dataset duration length and quality evaluation score.
Life-log Category Max (min) Mean (min) Min (min)
Short 580 142 23
Medium 760 319 118
Long 984 516 220
@@
Our constructed video selection library consisting of 7852 videos, covers a wide range of daily life scenarios and scenes across different times of the day, as shown in Fig. 3 (a), (c) and (d), a total of 135 scenarios are covered, here we show the top 30 most frequent occurrences in Fig. 3 (a). The statistics of the time period and scene information are shown in Fig. 3 (c), and (d) respectively. In addition, due to the commonality of human activities (e.g., eating, cooking), certain videos in the video selection library are selected multiple times. As shown in Fig. 3 (f), the vast majority of videos in X-LeBench appear less than 5 times, with a small number of videos appearing more than 5 times, reflecting the diversity of our simulations while preserving the representativeness of repeated contexts.
Table 3: The number of annotations across different tasks.
@@
3.2 Benchmark Tasks
The goal is to comprehensively evaluate the extremely long video understanding capabilities of current systems and models. As shown in Tab. 4, X-LeBench introduces a suite of evaluation tasks specifically designed for daily-life long-form egocentric videos. These tasks encompass object-, people-, and moment-related temporal localization, multi-level summarization, action counting, and summary ordering.
@@ 3.2 Benchmark Tasks
Table 4: Tasks in this benchmark and their corresponding examples. Q and A denote query and answer examples, respectively.

### v2

@@ Abstract
Long-form egocentric video understanding provides rich contextual information and unique insights into long-term human behaviors, holding significant potential for applications in embodied intelligence, long-term activity analysis, and personalized assistive technologies. However, existing benchmark datasets primarily focus on single, short (e.g., minutes to tens of minutes) to moderately long videos, leaving a substantial gap in evaluating extensive, ultra-long egocentric video recordings. To address this, we introduce X-LeBench, a novel benchmark dataset meticulously designed to fill this gap by focusing on tasks requiring a comprehensive understanding of extremely long egocentric video recordings. Our X-LeBench develops a life-logging simulation pipeline that produces realistic, coherent daily plans aligned with real-world video data. This approach enables the flexible integration of synthetic daily plans with real-world footage from Ego4D—a massive-scale egocentric video dataset covers a wide range of daily life scenarios—resulting in 432 simulated video life logs spanning from 23 minutes to 16.4 hours. The evaluations of several baseline systems and multi-modal large language models (MLLMs) reveal their poor performance across the board, highlighting the inherent challenges of long-form egocentric video understanding, such as temporal localization and reasoning, context aggregation, and memory retention, and underscoring the need for more advanced models.
X-LeBench: A Benchmark for Extremely Long Egocentric Video Understanding
[…]
@@ 1 Introduction
Understanding long-form egocentric videos captured from a first-person perspective over extended periods holds significant potential for advancing various domains such as embodied intelligence, long-term activity analysis, and personalized assistive technologies (plizzari2024outlook; lv2024aria; park2016egocentric). These videos provide rich contextual information and unique insights into human behaviors as they unfold naturally throughout the day. The ability to analyze such extensive recordings is key to developing more personalized agent systems that can construct long-term memory, anticipate user needs, and interact seamlessly in real-world settings lin2022egocentric; jia2022egotaskqa; pramanick2023egovlpv2.
@@
However, most existing datasets soomro2012ucf101; caba2015activitynet; kay2017kinetics; li2020hero; xiao2021next; miech2019howto100m, predominantly feature short individual video clips captured from third-person views, making them insufficient for understanding the continuous, nuanced context of daily human activities or for in-depth human-centered research that requires a first-person perspective. While more recent benchmarks and datasets focus on egocentric video understanding  damen2018scaling; damen2022rescaling; grauman2022ego4d; zhu2023egoobjects; mangalam2023egoschema; grauman2024ego; lv2024aria; fan2025benchmarks, they remain limited in capturing continuous, long-term human activities. For instance, Ego4D grauman2022ego4d features massive-scale egocentric videos covering a wide range of daily activities, with video durations ranging from 5 seconds to 7 hours, yet subsequent research mangalam2023egoschema; grauman2024ego; islam2024video; rodin2024action; chandrasegaran2024hourvideo often focuses on isolated clips or recordings, falling short of capturing the full scope of long-term daily human activities. This limitation hinders the evaluation of models designed to process ultra-long video streams. In particular, it restricts the ability to challenge current long-form video understanding systems and models that construct long-term memory from video recordings and retrieve relevant information in response to user queries. Without extensive, continuous contextual data, evaluating the robustness of their performance becomes difficult.
@@
Creating benchmark datasets that span several hours of egocentric video is necessary but presents significant challenges. Data acquisition is a primary hurdle, requiring participants to wear recording devices for extended periods is labor-intensive and raises privacy concerns. Device limitations, including storage constraints and reliability issues, further complicate continuous video capture. Additionally, annotating long-form videos is time-consuming and prone to annotator fatigue, affecting label accuracy and consistency.
@@
To address these challenges, we introduce X-LeBench, a versatile and scalable benchmark dataset designed for evaluating tasks on extremely long egocentric videos. X-LeBench features a life-logging simulation pipeline that simulates extended video logs by integrating short (seconds) or moderately long (hours) video clips with dynamically generated daily plans. Leveraging large language models (LLMs), it generates realistic, contextually rich schedules aligned with real-world activities based on adjustable input settings. Specifically, this simulation integrates synthetic daily plans with actual footage from Ego4D, then iteratively optimizes the simulation process based on retrieved information, producing video life logs that mirror daily activities in rich contexts with duration extended to dozens of hours. A generated sample is shown in Fig. 1, showcasing video segments with different timestamps. Notably, X-LeBench offers a customizable and scalable design, enabling the synthesis of datasets with various durations and content to accommodate diverse research needs.
@@
Our initial evaluations of baseline systems and multi-modal LLM (MLLM) reveal consistently poor performance on X-LeBench, highlighting the inherent difficulties of long-form egocentric video understanding and underscoring the need for more advanced models capable of interpreting and analyzing ultra-long egocentric videos.
@@
Our contributions are: (1) We present X-LeBench, the first benchmark dataset that encompasses ultra-long egocentric video recordings; (2) We introduce a novel and customizable pipeline that simulates realistic, hours-long egocentric video life logs by integrating synthetic daily plans with real-world footage; (3) We conduct extensive evaluations of existing models on X-LeBench, exposing significant performance gaps and key challenges for future research.
@@ 2 Related Works
2.1 Egocentric Video Benchmarks
Egocentric video understanding has received increasing attention cheng2024egothink; huang2024vinci. Ego4D is a landmark dataset, offering 3,670+ hours of egocentric footage across 74 locations, facilitating significant progress in the field.
Related works bain2022whisperx; kevin2022egovlp; pramanick2023egovlpv2; islam2024video extend Ego4D, further exploring various applications of egocentric video understanding. EgoSchema mangalam2023egoschema offers an egocentric video question-answering benchmark with over 5,000 curated multiple-choice pairs but focuses on 3-minute clips. The AEA dataset lv2024aria offers multi-modal egocentric data, HourVideo chandrasegaran2024hourvideo curates hour-long videos from Ego4D for evaluating video-language understanding. EgoPlan-Bench2 qiu2024egoplan assesses planning capabilities of MLLMs. However, these works either lack coherent, continuous, long-term daily-life recordings or face data scale and diversity limitations.
Recently, EgoLife’s week-long recordings yang2025egolife offer valuable long-context data, but high costs, operational issues, limited 6 subjects, and an indoor focus restrict its scalability and diversity. In contrast, we developed a novel, cost-effective pipeline to synthesize ultra-long, coherent egocentric video life logs from existing datasets, offering a scalable and extensible alternative and significantly broadening the scope and applicability of long-term egocentric video understanding.
@@ 2.1 Egocentric Video Benchmarks
Dataset
Avg. Duration
Annotation
Scheme
Egocentric #QAs #Data
MVBench 0.27 Auto No 4,000 3,641
ActivityNet-QA 1.85 Manual No 8,000 800
EgoSchema 3 Auto&Manual Yes 5,063 5,063
EgoPlan-Bench2 up to 5 Auto&Manual Yes 1,321 1,113
MovieChat-1K 9.4 Manual No 1,950 130
MLVU 12 Auto&Manual Partial 2,593 757
Video-MME (Short) 1.37 Manual No 2,700 900
@@
HourVideo 45.7 Auto&Manual Yes 12,976 500
InfiniBench 76.34 Auto&Manual No 108,200 1219
LVBench 68.35 Manual No 1,594 103
EgoLife 2,658 Auto&Manual Yes 3,000 6
Ours (Short) 142 Auto&Manual Yes 26,932 432
@@
2.2 Long-form Video Benchmarks
The definition of “long” in video understanding varies across benchmarks, with durations ranging from minutes to hours. As shown in Tab. 1, EgoSchema defines 3-minute videos as long using their proposed certificate length, Video-MME fu2024video classifies videos with a length of 30-60 minutes as long. HourVideo and LVBench wang2024lvbench define long videos as 20+ and 30+ minutes, respectively. MLVU zhou2024mlvu offers videos of diversified lengths with a 12-minute average length, providing comprehensive evaluation tasks for MLLMs’ long video understanding capabilities. InfiniBench ataallah2024infinibench pushes the boundary with 50-minute videos and over 108,000 question-answer pairs, posing significant challenges for leading AI models.
Despite recent advances, existing datasets remain insufficiently long for evaluating ultra-long video processing and rarely focus on egocentric content. X-LeBench fills this gap by redefining “long video” to include multi-hour, contextually consistent egocentric recordings with rich annotations to advance model and system development.
@@ 2.3 LLM-assisted Annotation Scheme
Traditional annotation processes require extensive human effort, but advances in LLMs and MLLMs achiam2023gpt; yao2024minicpmvgpt4vlevelmllm; Qwen-VL; team2023gemini; touvron2023llama; lin2023video have facilitated automated annotation in benchmarks. As shown in Tab. 1, several video understanding benchmarks leverage LLMs/MLLMs to streamline annotation process mangalam2023egoschema; rawal2024cinepile; patraucean2023perception; li2024mvbench. For instance, EgoSchema generates question-answer pairs by querying an LLM with human-annotated narrations and tailored prompts.
Given the multi-hour duration of our dataset, manual annotation is impractical and prone to fatigue-induced errors. To address this, we also adopt an automatic and manual annotation scheme. First, we adapt Ego4D’s manual annotations to align with our task requirements. Next, we use LLMs to consolidate summaries of 5-minute video clips from Ego4D at various granularities, creating single-video, multi-video, and holistic-level summaries (detailed in Appendix C).
@@ 3 X-LeBench
In real-world scenarios, continuous egocentric life-logging is often constrained by hardware limitations and privacy concerns, making uninterrupted long-term recordings infeasible. Consequently, extended life logs must be reconstructed from multiple video segments captured at different times. Existing benchmark datasets lack the structural continuity and task setup required for studying temporally extended, cross-hour behaviors and reasoning.
Building on this insight, we develop the life-logging simulation pipeline, enabling scalable and customizable simulation of hours-long video life logs, while maintaining reality and contextual coherence. We also define tasks that better reflect real-world long-term requirements and go beyond Ego4D’s original scope. The following sections detail our methodology for X-LeBench creation.
@@ 3.1 Life-logging Simulation Pipeline
Stage 1 - Persona Generation. To ensure the simulations better reflects the multifaceted nature of real-world human activity, enriching the realism and variability of the resulting dataset, we dynamically generate personalized character profiles based on different predefined locations and the Myers-Briggs Type Indicators (MBTI) myers1962myers, generating basic background information including the character’s personality traits, lifestyle, hobbies and general daily routines. By incorporating varied character settings, we capture a wide range of behavioral variations that enrich our dataset. This pragmatic choice enables the generation of diverse daily plans and activity patterns, promoting variability across simulations.
Each daily plan is then segmented into time-specific activity chunks, ensuring structured activity distribution. During this stage, we use GPT-4o hurst2024gpt for character generation, setting 9 different locations based on Ego4D’s demographic data, and 16 MBTI types per location, resulting in 144 diverse persona profiles (profile example is shown in Appendix A).
@@
Stage 2 - Video Information Extraction. We construct our video selection library by carefully selecting 7852 videos labeled with scenario information and dense clip-level summaries from the Ego4D, excluding redacted content. Then we extract three key attributes from each video, including the time, scene, and the main content. Specifically, we use Gemini-1.5-Pro team2024gemini to analyze videos and transform information. To facilitate simulation with reasonable environment settings and lighting conditions, videos are categorized into different scene environments: indoor, outdoor, and mixed, as well as different time periods: daytime, nighttime, twilight, and uncertain (e.g., indoor videos lacking external cues are labeled “uncertain”). Content information is derived from Ego4D scenario tags, and 1-sentence summaries consolidated from 5-minute clip-level summaries from each video. This structured metadata ensures precise alignment between video content and the generated daily plans in the next stage.
@@
Figure 3: Dataset statistics. (a) Video selection library consists of 7852 videos from Ego4D dataset, covering 135 scenarios, here shows the top 30. (b) X-LeBench has video life-log durations spanning from 23 minutes to 16.4 hours, here shows distribution of duration range across different data categories. (c) and (d) show the distribution of scene and time information categories of videos in video selection library, respectively. (e) Statistical information on duration lengths of different data categories, including minimum, mean, median and maximum values. (f) The distribution of number of videos with different occurrences.
@@
Stage 3 - New Dataset Generation. In this stage, we match the generated daily plan chunks (Stage 1) with videos (Stage 2) to construct coherent video life-log simulations. For each plan chunk, we retrieve a video that aligns with its time, scene, and content, the selection process follows rule below:
$V={\underset{V\in\mathbf{M_{v}}}{{\arg\max}}\ S(\mathbf{M_{v}}(C_{t},C_{scene},C_{scenario}),C_{desc})}$
@@
where $V$ is the selected video for chunk $C$, determined by the alignment with $C$’s core information, which is time ($C_{t}$), scene ($C_{scene}$), scenario ($C_{scenario}$), and 1-sentence summary description ($C_{desc}$). $\mathbf{M_{v}(\cdot)}$ denotes the operation of matching videos based on the provided chunk information. $S(\cdot)$ represents the sentence similarity computation between $C_{desc}$ and 1-sentence summary set of matching videos. Specifically, $C_{t}$ and $C_{scene}$ are inferred from chunk timestamps and content, while $C_{scenario}$ is inferred using the dataset’s predefined scenario list (i.e., the dataset reference in Fig. 2).
@@
To efficiently identify the most suitable videos from the large-scale video selection library, selection process follows a coarse-to-fine retrieval strategy. We first apply a coarse-grained filtering step, i.e., the operation $\mathbf{M_{v}}(C_{t},C_{scene},C_{scenario})$, a rule-based matching to narrow down the candidate video set. Finally, to refine the selection, we calculate the sentence similarity between $C_{desc}$ and the summaries of the candidate video set $\mathbf{M_{v}}$, ensuring alignment across time, scene, and content dimensions.
@@
In addition, to enhance the contextual reasonableness and coherence of overall video life logs, we preferentially select videos recorded in the same location for each simulation. Furthermore, an iterative optimizing strategy is adopted, i.e., after each matching of a daily plan chunk, we update the life-log simulation memory with the information of selected video, require LLM uses this context to reasonably adjust and update the subsequent daily plan chunks, ensuring that the overall simulation maintains logical coherence and contextual alignment. Details can be found in Appendix B.
Statistics. Our life-logging simulation pipeline allows for flexible customization by inputting various locations and MBTI types, generating diverse persona profiles and simulations. We can customize different numbers of chunks to get video life logs of various video compositions. In this work, we set three different chunk numbers for simulation, namely 4, 9, and 15 (corresponding to short, medium, and long life-log categories) for each location-MBTI combination generated persona, thus obtaining 432 extremely long video life logs. To assess dataset quality, we compare our method with a randomly sampled baseline (Appendix E), our method retrieves videos with contextualized timestamps and more consistent content that closely resemble real-life recorded scenes. We also conduct a five-point human evaluation on realism and contextual consistency. As summarized in Tab. 2, Fig. 3 (b) and (e), our dataset achieved scores over 4 across all categories, demonstrating strong alignment with real-world activity patterns and logical continuity. Evaluation details are provided in the Appendix D. Also, the dataset exhibits diverse duration distributions: Short data are mostly 1-2 hours; Medium data are mostly 4-5 hours; And long data are mostly 6-8 hours. With the average duration of 2.37, 5.32 and 8.6 hours.
Table 2: Statistics of dataset duration length and quality.
Life-log Category
Max
Mean
Min
Realism
Contextual
Consistency
Short 580 142 23 4.71 4.50
Medium 760 319 118 4.26 4.37
Long 984 516 220 4.41 4.02
@@
Our video selection library contains 7852 videos, covering a wide range of daily life scenarios and scenes across different times of the day, a total of 135 scenarios are covered. Fig. 3 (a), (c) and (d) illustrate the top 30 most frequent scenarios, time distributions, and scene types. In addition, due to the commonality of human activities (e.g., eating, cooking), certain videos are selected multiple times. Fig. 3 (f) shows that most videos appear fewer than five times, with a small number of videos appearing more than 5 times, reflecting the diversity of our simulations while preserving the representativeness of repeated contexts.
Table 3: The number of annotations.
@@
Table 4: Tasks in this benchmark and their corresponding examples. Q and A denote query and answer examples.
3.2 Benchmark Tasks
As shown in Tab. 4, X-LeBench introduces a suite of evaluation tasks specifically designed for daily-life long-form egocentric videos. These tasks encompass object-, people-, and moment-related temporal localization, multi-level summarization, action counting, and summary ordering. Details and examples are in Appendix C.